OpenStudy (anonymous):

5.3(29 - 3.33x) 34.3 - 8.63x 153.7 - 17.649x 153.7x - 17.649 17.649 + 153.7x

TheSmartOne (thesmartone):

Do you know the distibutive property? @cjackson2002

OpenStudy (anonymous):

kinda!

TheSmartOne (thesmartone):

\(\Huge a(b-c)=(a \times b)-(a \times c)\)

OpenStudy (anonymous):

yea?

TheSmartOne (thesmartone):

Use that.

OpenStudy (anonymous):

uhhhhhh???? nvm!

TheSmartOne (thesmartone):

Are you getting confused because of the x?

OpenStudy (anonymous):

no

TheSmartOne (thesmartone):

Did you get an answer?